Given the wave equation y(x,t) =2.0 Sin (3.0x-4.0t + π/2)m, where x is in meter and t is in second. Find the:

a) amplitude

b)Frequency

c) wavelength

d)speed

e) initial height at x=1.0m


1
Expert's answer
2021-01-31T03:40:04-0500

The general equation of wave is:


"y(x,t) = A\sin(kx-\omega t + \varphi_0)"

where "k = \dfrac{2\pi}{\lambda}" is the wave number ("\lambda" is the wavelength), "\omega = 2\pi f" is the angular frequency ("f" is the frequency), "A" is the amlitude, and "\varphi_0" is the initial phase.

Comparing the general equation with the given one, find:


"A = 2m""\omega = 4rad/s \Rightarrow f =\dfrac{4}{2\pi} \approx 0.64Hz""k = 3m^{-1}\Rightarrow\lambda = \dfrac{2\pi}{3} \approx 2.09m"

Speed of the wave is given as:


"v = f\lambda = 0.64\cdot 2.06m/s \approx 1.33m/s"


Substituting "x = 1m", "t = 0", find:


"y(1,0) = 2\sin(3\cdot 1 - 4\cdot0 + \pi/2) \approx 0.16m"

Answer. a) 2m, b) 0.64Hz, c) 2.09m, d) 1.33 m/s, e) 0.16m